How many pounds of the $7.28 coffee should be mixed with 7 pounds of $6.42 coffee to sell the mixture for $7.20 per pound?
Answer by checkley77(12844)

7.28x+7*6.42=7.20(7+x)
7.28x+44.94=50.4+7.20x
7.28x-7.20x=50.4-44.94
.08x=5.46
x=5.46/.08
x=68.25 pounds of $7.28 coffee is used.
Proof:
7.28*68.25+44.94=7.20(7+68.25)
496.86+44.94=7.20*75.25
541.8=541.8